Hello how are you I was working on this problem n-29=-53and i wanted to get an understanding on working the problem.......this is what i did 53-29=24 and then i add the neg sign b/c both numbers were neg. so my answer was n=-24
Answer by stanbon(75887)

n-29=-53
----
Add 29 to both sides to solve for "n":
n-29+29 = -53+29
n = -24
===========
Cheers,
Stan H.
===========